OpenStudy (anonymous):

If L is the line through the point (- 3, 5) having a slope of , complete the equation of L . y =

OpenStudy (anonymous):

do u have any idea on this at all?

OpenStudy (anonymous):

@Sarahbelle

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Sarahbelle If L is the line through the point (- 3, 5) having a slope of , complete the equation of L . y = \(\color{blue}{\text{End of Quote}}\) having a slope of what ?

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \(\color{blue}{\text{Originally Posted by}}\) @Sarahbelle If L is the line through the point (- 3, 5) having a slope of , complete the equation of L . y = \(\color{blue}{\text{End of Quote}}\) having a slope of what ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Honestly, no. I'm using Odyssey and its terrible at explanations, and when I ask me teacher, she says she doesn't know.

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

so uh having the slope of?

Nnesha (nnesha):

oaky then there should be two order pair or graph or they should give u da slope

OpenStudy (anonymous):

i think when u pasted it the slope disappeared

OpenStudy (anonymous):

2/3

Nnesha (nnesha):

okay so y = mx + b is slope intercept form where m is slope and b is y-intercept so replace m by 2/3

OpenStudy (anonymous):

right... and plug the points into the equation....

OpenStudy (anonymous):

Okay, I did that.

Nnesha (nnesha):

okay so you have one order pair where first number is x and 2nd one is y (x , y ) so replace x and y by their values then solve for b

OpenStudy (anonymous):

B=22/3?

Nnesha (nnesha):

nope b is not 22/3

Nnesha (nnesha):

\[\huge\rm y = \frac{ 2 }{ 3 }x + b \] substitute x and y values

OpenStudy (anonymous):

b=7?

Nnesha (nnesha):

yep right :-) (-:

OpenStudy (anonymous):

Thanks for being so patient! Math is not my best skill XD

Nnesha (nnesha):

my pleasure :) and don't worry just practice then u gonna be expert :-)